# Foundational Automatic Evaluators: Scaling Multi-Task Generative Evaluator Training for Reasoning-Centric Domains

### v1

@@ 1 Introduction
The past two years has seen the rapid adoption of large language models (LLMs) as automatic evaluators in response to demands for scalable evaluation of LLM outputs. LLM-based evaluators serve as judges for popular benchmarks (Dubois et al., 2024; Zheng et al., 2023), generative reward models for preference optimization (Yuan et al., 2024; Wu et al., 2024), and verifiers/critics in inference-time scaling settings (McAleese et al., 2024; Zhou et al., 2025b). The widespread integration of evaluators into nearly every phase of the LLM development cycle (Wu, 2025) demands evaluators that can handle multiple evaluation tasks while operating effectively across diverse domains.
Different settings require different evaluation abilities: Alignment needs evaluators capable of comparing different responses, i.e., pairwise evaluation, whereas monitoring model outputs requires finding minute mistakes, i.e., step-level evaluation. More recently, generative evaluators are tasked with providing reward signals during reinforcement learning (RL) training (Team Kimi et al., 2025; Jiang et al., 2025b), and are expected to grow in importance as RL moves towards unverifiable domains (Gunjal et al., 2025; Jayalath et al., 2025) in complex reasoning settings (Ke et al., 2025a; Ferrag et al., 2025). As evaluators take central roles in training and evaluating the next generation of models, they must be flexible enough to evaluate as the setting demands.
Compounding the challenges of multi-task evaluation is the expanding number of domains requiring evaluation: RL-training has quickly moved from math reasoning (e.g., Yu et al. (2025b)) to general-purpose reasoning (Ma et al., 2025) (e.g., history or economics). Agentic settings introduce additional wrinkles: With autonomously acting single agents (OpenAI, 2025; Nguyen et al., 2025; Wei et al., 2025) and complex multi-agent workflows (Liang et al., 2025; Alzubi et al., 2025) now being set free to browse the web and act on behalf of users with minimal oversight, evaluators must assess not only agent reasoning, but also proposed tool-use. These systems, sometimes built with intricate model-generated (Hu et al., 2024b; Zhang et al., 2024a; Ke et al., 2025b) interdependencies, are bottlenecked, in part, by subpar evaluation (Cemri et al., 2025).
@@ 2 Background and Related Work
Past work in generative automatic evaluators. Capable LLMs, like GPT-4, were originally prompted as scalable evaluators (Wang et al., 2023a; Liu et al., 2023b; Fu et al., 2024; Chiang & Lee, 2023). Subsequent analysis revealed pitfalls of prompted approaches, like biases with respect to position (Wang et al., 2023b; Li et al., 2023), length (Zeng et al., 2023; Park et al., 2024), or self-preference (Panickssery et al., 2024). Finetuning specialized evaluators emerged as a result, with early approaches using teacher model outputs to do supervised finetuning (SFT) (Kim et al., 2023; 2024b; Li et al., 2023; Park et al., 2024; Shiwen et al., 2024) or direct preference optimization (DPO) (Hu et al., 2024c; Ye et al., 2024), often focusing only on one or two evaluation tasks. More recent methods moved to reasoning models as teachers (Khalifa et al., 2025).
@@
Vu et al. (2024); Wang et al. (2024a); Cao et al. (2024); Alexandru et al. (2025) train foundational evaluators at larger data scales with multi-protocol capabilities via offline training methods like SFT or DPO. Such approaches take inspiration from general-purpose, large-scale multi-task learning (Sanh et al., 2021; Raffel et al., 2020; Wei et al., 2021), which showed broad generalization capabilities emerge with the scaling of training data. Foundational evaluators likewise were empirically shown to generalize to unseen evaluation tasks, prompts, and criteria while being more robust to common biases (Vu et al., 2024; Wang et al., 2024a).
@@
Recent work has focused on methodological advances, either using inference-time scaling (Liu et al., 2025d; Chan et al., 2025; Zhao et al., 2025) or online training like reinforcement learning from verifiable rewards (RLVR) (Chen et al., 2025a; b; Whitehouse et al., 2025; Xu et al., 2025b; Xiong et al., 2025b) to improve evaluator performance. Because RLVR is computationally demanding with relatively brittle training pipelines (Guo et al., 2025; Yang et al., 2025), recent evaluators are typically trained on a small amount of data for a single task. Our work bridges early work in training foundational evaluators with more recent methodological advancements, demonstrating that a simple semi-online training approach enables stable multi-task training at scale.
@@
Desiderata for a new generation of evaluators. Here, we outline our design philosophy for FARE. Beyond accuracy and robustness, we seek efficiency, as many evaluation settings like inference-time reranking or RL rollout verification demand low latency. In contrast to recent long chain-of-thought (CoT) evaluators (Chen et al., 2025b; Khalifa et al., 2025), we select base models with either no or very compact “thinking” CoTs. We also explicitly avoid having the evaluator generate reference answers. Past work has used evaluators to generate references during evaluation (Zheng et al., 2023; Li et al., 2024) or training rollout (Chen et al., 2025b). Not only does this risk severely degrading performance when the reference is wrong (Krumdick et al., 2025), it also converts the relatively easier evaluation task into a harder generation task (Zhou et al., 2025a).
@@ 3.1 Data curation
Existing data consists of training samples from proven sources that have produced effective evaluators (Vu et al., 2024; Cao et al., 2024). These datasets consist of high quality annotations from humans and frontier LLMs, and cover evaluation tasks like step-level, single-rating, and pairwise evaluation and domains like chat quality, code, and safety. Following Wang et al. (2024a), we largely focus our data collection on modern (2024 and beyond) datasets, as these datasets contain the most up-to-date model responses and fresh annotations. Beyond evaluator-specific training, we take advantage of existing preference fine-tuning datasets used for RLHF (Ouyang et al., 2022) and DPO training (Rafailov et al., 2023), converting these directly into pairwise evaluation samples. In domains with objectively correct answers, e.g., math, we also create verification training data with positive responses as correct/reference responses and negative responses as incorrect responses.
@@
  • Generate-then-grade. Here, we leverage a mix of recent and established training datasets comprised of question $q$ and verifiable ground-truth answer $a$. We sample up to $20$ responses per $q$ from various generator models, then grade each responses based on $a$. After grouping responses by correctness, we create verification and pairwise samples. We use 12 unique generators from 6 model families, covering reasoning and non-reasoning models. This ensures that FARE are trained on a diverse array of model responses, enabling better generalization across distinct response distributions (Singh et al., 2025). Generate-then-grade enables us to incorporate problems from recent, challenging datasets covering frontier math and reasoning tasks. This enhances the quality of our pairwise and verification data with difficult-to-evaluate reasoning-focused samples.
@@ 3.2 Model training
These established approaches have their limitations: Teacher models introduce distribution shifts with respect to the policy model, a common problem with imitation learning (Ross et al., 2011), with choice of teacher model having a large impact on downstream performance (Guha et al., 2025). On the other hand, RL training is compute and time-intensive, making it difficult to scale to large data quantities, with past work (Liu et al., 2025c) only exploring small (1.5B parameter) models via ad-hoc interventions like reference policy resetting.
We borrow desirable qualities from both paradigms and use semi-online iterative rejection sampling SFT (RS-SFT) (Touvron et al., 2023; Dong et al., 2023), which was recently shown to be competitive with RLVR approaches (Xiong et al., 2025a). RS-SFT avoids sub-optimal distribution shift by finetuning on correct evaluation traces produced by the policy model while employing a computationally lightweight policy update step. This enables simple yet stable scaling to millions of training samples without sampling from a teacher model. An overview of our approach is shown in Fig. 1.
@@
Our approach draws inspiration from algorithms such as STaR (Zelikman et al., 2022) and RAFT (Dong et al., 2023), with some key differences. STaR notably re-initializes training from $\pi_{\theta_{0}}$ for each iteration $t$ and samples only one greedy response per input prompt, while RAFT relies on an external reward model to rank generated outputs. Because the automatic evaluation setting is inherently verifiable, we omit the need for a reward model to rank sampled responses.
Specific to evaluators, the Self-Taught Evaluator (STE) paradigm of Wang et al. (2024b) and follow-up EvalPlanner (Saha et al., 2025) are closely related to our approach. STE follows STaR with policy re-initialization and EvalPlanner uses multiple SFT and DPO training runs per iteration $t$. Further, these works use in-the-loop synthetic data generation, sampling responses to a small number (<25<25K) of seed questions from a fixed generator. Pairwise samples are then created from correct/incorrect responses and used to train the model. This data generation approach cannot be adapted to create data for other tasks, like step-level evaluation, fundamentally limiting the task abilities of STE. A secondary concern is a lack of exposure to diverse response distributions; Evaluations in Wang et al. (2024a) show that scaling training data with a simpler training recipe leads to better generalization across benchmarks compared to STE.
@@
Inclusion of direct judgment data. Past work (Wang et al., 2024a; Cao et al., 2024) has showed the importance of including direct judgment data samples to isolate judgment training signal. These are samples where the critique $c$ is omitted, and the input protocol $p$ is modified to prompt directly for a judgment. To precisely control the fraction of direct judgment data, we convert a fixed percent of $\mathcal{D}_{t}$ to direct judgment data by dropping generated critiques and modifying the input prompt accordingly. In App. D, we ablate the proportion of direct judgment data and show such data enables FARE to be prompted to exclude critiques for faster inference.
@@
Base models. We train two models starting from Qwen3-8B-Base (Yang et al., 2025) and gpt-oss-20B (Agarwal et al., 2025), denoted FARE-8B and FARE-20B. We find Qwen3-8B (post-trained) to be over-trained, and therefore cold-start Qwen3-8B-Base from SFT data from Qwen2.5-32B-Instruct, which we denote Qwen3-8B-ColdStart. See § B.2 for additional details.
@@ 4.1 Core benchmarks
Setup. We evaluate along five diverse aspects: (ireasoning with JudgeBench (Tan et al., 2024), ReasoningJudgeBench (RJB) (Xu et al., 2025b), and PPE Correctness (Frick et al., 2024), (iibias and robustness with RM-Bench (Liu et al., 2024b), (iiitool-use with When2Call (Ross et al., 2025), (ivstep-level error identification with ProcessBench (Zheng et al., 2024), and (vreference-based verification with VerifyBench (Yan et al., 2025).
For RM-Bench and PPE Correctness pairwise benchmarks, we adopt default evaluation setups, running each benchmark once with a fixed random ordering of responses. For other pairwise benchmarks, we report consistent-accuracy (Tan et al., 2024), where each test sample is run twice, swapping the order of response A and response B. If the evaluator selects a different response between runs (i.e., is positionally biased), then the sample is marked incorrect; if the evaluator is consistent, the judgment is graded against the ground-truth. For ProcessBench and VerifyBench, we report F1-score [^1] ProcessBench defines their reported F1-score differently from the traditional F1-score; See this link. and accuracy, respectively. We compare FARE against other finetuned generative and prompted evaluators. We report official numbers from past benchmarks, reporting sources in App. C. If necessary, we run each baseline using its own prompt template. For ProcessBench, we additionally compare against non-generative process reward models (PRMs).
@@
FARE also are extremely strong step-level evaluators, as shown by ProcessBench performance in Table 2. FARE-8B is the best small-sized generative critic model, outperforming the recently released StepWiser-7B (Xiong et al., 2025b), a RL-trained specialized step-level evaluator, by 1.6 points. FARE-20B outperforms the specialized Qwen2.5-Math-72B-PRM by 6.1 points, even matching GPT-5. Most notably, FARE excel on the two most challenging splits, OlympiadBench and OmniMATH, with FARE-8B beating StepWiser-7B by 5.6 points and FARE-20B beating Qwen2.5-Math-72B-PRM by 7.3 points on average. Overall, FARE are not only capable outcome-level evaluators, but are able to find subtle mistakes that manifest at the step-level.
FARE are also capable verifiers, as shown in Table 3, with both models outperforming all reported baselines on VerifyBench-Hard (Yan et al., 2025). In particular, FARE-20B excels beats the next best model, GPT-4o, by 12.5 absolute points. As we demonstrate in § 4.2, when used as verifiers during GRPO settings, FARE bring tangible benefits over typical string-matching verifiers.
Scaling inference-time compute. Recently, sampling parallel judgments and aggregating via majority vote, i.e., self-consistency (Wang et al., 2022), has been used to improve evaluator performance. In Table 4, we use self-consistency with 32 responses (SC@32) on PPE, comparing with J1 and DeepSeek-GRM (Liu et al., 2025d). DeepSeek-GRM also trains a MetaRM to perform judgment re-ranking at test-time. Across most splits, using SC@32 improves performance, with up to a 3.3 point improvement for FARE-20B on the MMLU-Pro split. Even without SC@32, FARE-8B beats DeepSeek-GRM-27B + MetaRM, with the gap widening with extra compute. Similarly, the gap between FARE-20B and J1-70B grows with extra compute. Interestingly, we see that across both our models, performance on MBPP+ slightly degrades, indicating that SC may not be the optimal way to use compute across all domains. Nonetheless, we observe gains in the aggregate.

### v2

@@ 1 Introduction
The past two years has seen the rapid adoption of large language models (LLMs) as automatic evaluators in response to demands for scalable evaluation of LLM outputs. LLM-based evaluators serve as judges for popular benchmarks (dubois2024length; zheng2023judging), generative reward models for preference optimization (yuan2024self; wu2024meta), and verifiers/critics in inference-time scaling settings (mcaleese2024llm; zhou2025evaluating). The widespread integration of evaluators into nearly every phase of the LLM development cycle (wu2025sailing) demands evaluators that can handle multiple evaluation tasks while operating effectively across diverse domains.
Different settings require different evaluation abilities: Alignment needs evaluators capable of comparing different responses, i.e., pairwise evaluation, whereas monitoring model outputs requires finding minute mistakes, i.e., step-level evaluation. More recently, generative evaluators are tasked with providing reward signals during reinforcement learning (RL) training (team2025kimi; jiang2025pag), and are expected to grow in importance as RL moves towards unverifiable domains (gunjal2025rubrics; jayalath2025compute) in complex reasoning settings (ke2025survey; ferrag2025llm). As evaluators take central roles in training and evaluating the next generation of models, they must be flexible enough to evaluate as the setting demands.
Compounding the challenges of multi-task evaluation is the expanding number of domains requiring evaluation: RL-training has quickly moved from math reasoning (e.g., yu2025dapo) to general-purpose reasoning (ma2025general) (e.g., history or economics). Agentic settings introduce additional wrinkles: With autonomously acting single agents (openai2025deepresearch; nguyen2025sfr; wei2025webagent) and complex multi-agent workflows (openmanus2025; alzubi2025opendeepsearch) now being set free to browse the web and act on behalf of users with minimal oversight, evaluators must assess not only agent reasoning, but also proposed tool-use. These systems, sometimes built with intricate model-generated (hu2024automated; zhang2024aflow; ke2025mas) interdependencies, are bottlenecked, in part, by subpar evaluation (cemri2025multi).
@@ 2 Background and Related Work
Past work in generative automatic evaluators. Capable LLMs, like GPT-4, were originally prompted as scalable evaluators (wang2023chatgpt; liu2023g; fu2024gptscore; chiang2023can). Subsequent analysis revealed pitfalls of prompted approaches, like biases with respect to position (wang2023large; li2023generative), length (zeng2023evaluating; park2024offsetbias), or self-preference (panickssery2024llm). Finetuning specialized evaluators emerged as a result, with early approaches using teacher model outputs to do supervised finetuning (SFT) (kim2023prometheus; kim2024prometheus; li2023generative; park2024offsetbias; skyworkcritic2024) or direct preference optimization (DPO) (hu2024themis; ye2024beyond), often focusing only on one or two evaluation tasks. More recent methods moved to reasoning models as teachers (khalifa2025process).
@@
vu2024foundational; wang2024direct; cao2024compassjudger; alexandru2025atla train foundational evaluators at larger data scales with multi-protocol capabilities via offline training methods like SFT or DPO. Such approaches take inspiration from general-purpose, large-scale multi-task learning (sanh2021multitask; raffel2020exploring; wei2021finetuned), which showed broad generalization capabilities emerge with the scaling of training data. Foundational evaluators likewise were empirically shown to generalize to unseen evaluation tasks, prompts, and criteria while being more robust to common biases (vu2024foundational; wang2024direct).
@@
Recent work has focused on methodological advances, either using inference-time scaling (liu2025inference; chan2025j1; zhao2025genprm) or online training like reinforcement learning from verifiable rewards (RLVR) (chen2025judgelrm; chen2025rmr1rewardmodelingreasoning; whitehouse2025j1; xu2025j4r; xiong2025stepwiser) to improve evaluator performance. Because RLVR is computationally demanding with relatively brittle training pipelines (dsr1_guo2025deepseek; yang2025qwen3), recent evaluators are typically trained on a small amount of data for a single task. Our work bridges early work in training foundational evaluators with more recent methodological advancements, demonstrating that a simple semi-online training approach enables stable multi-task training at scale.
@@
Desiderata for a new generation of evaluators. Here, we outline our design philosophy for FARE. Beyond accuracy and robustness, we seek efficiency, as many evaluation settings like inference-time reranking or RL rollout verification demand low latency. In contrast to recent long chain-of-thought (CoT) evaluators (chen2025rmr1rewardmodelingreasoning; khalifa2025process), we select base models with either no or very compact “thinking” CoTs. We also explicitly avoid having the evaluator generate reference answers. Past work has used evaluators to generate references during evaluation (zheng2023judging; li2024crowdsourced) or training rollout (chen2025rmr1rewardmodelingreasoning). Not only does this risk severely degrading performance when the reference is wrong (krumdick2025no), it also converts the relatively easier evaluation task into a harder generation task (zhou2025variation).
@@ 3.1 Data curation
Existing data consists of training samples from proven sources that have produced effective evaluators (vu2024foundational; cao2024compassjudger). These datasets consist of high quality annotations from humans and frontier LLMs, and cover evaluation tasks like step-level, single-rating, and pairwise evaluation and domains like chat quality, code, and safety. Following wang2024direct, we largely focus our data collection on modern (2024 and beyond) datasets, as these datasets contain the most up-to-date model responses and fresh annotations. Beyond evaluator-specific training, we take advantage of existing preference fine-tuning datasets used for RLHF (instructgpt_ouyang2022training) and DPO training (dpo_rafailov2023direct), converting these directly into pairwise evaluation samples. In domains with objectively correct answers, e.g., math, we also create verification training data with positive responses as correct/reference responses and negative responses as incorrect responses.
@@
  • Generate-then-grade. Here, we leverage a mix of recent and established training datasets comprised of question $q$ and verifiable ground-truth answer $a$. We sample up to $20$ responses per $q$ from various generator models, then grade each responses based on $a$. After grouping responses by correctness, we create verification and pairwise samples. We use 12 unique generators from 6 model families, covering reasoning and non-reasoning models. This ensures that FARE are trained on a diverse array of model responses, enabling better generalization across distinct response distributions (singh2025shelf). Generate-then-grade enables us to incorporate problems from recent, challenging datasets covering frontier math and reasoning tasks. This enhances the quality of our pairwise and verification data with difficult-to-evaluate reasoning-focused samples.
@@ 3.2 Model training
These established approaches have their limitations: Teacher models introduce distribution shifts with respect to the policy model, a common problem with imitation learning (ross2011reduction), with choice of teacher model having a large impact on downstream performance (guha2025openthoughts). On the other hand, RL training is compute and time-intensive, making it difficult to scale to large data quantities, with past work (liu2025prorl) only exploring small (1.5B parameter) models via ad-hoc interventions like reference policy resetting.
We borrow desirable qualities from both paradigms and use semi-online iterative rejection sampling SFT (RS-SFT) (touvron2023llama; dong2023raft), which was recently shown to be competitive with RLVR approaches (xiong2025minimalist). RS-SFT avoids sub-optimal distribution shift by finetuning on correct evaluation traces produced by the policy model while employing a computationally lightweight policy update step. This enables simple yet stable scaling to millions of training samples without sampling from a teacher model. An overview of our approach is shown in Fig. 1.
@@
Our approach draws inspiration from algorithms such as STaR (zelikman2022star) and RAFT (dong2023raft), with some key differences. STaR notably re-initializes training from $\pi_{\theta_{0}}$ for each iteration $t$ and samples only one greedy response per input prompt, while RAFT relies on an external reward model to rank generated outputs. Because the automatic evaluation setting is inherently verifiable, we omit the need for a reward model to rank sampled responses.
Specific to evaluators, the Self-Taught Evaluator (STE) paradigm of wang2024self and follow-up EvalPlanner (saha2025learning) are closely related to our approach. STE follows STaR with policy re-initialization and EvalPlanner uses multiple SFT and DPO training runs per iteration $t$. Further, these works use in-the-loop synthetic data generation, sampling responses to a small number (<25<25K) of seed questions from a fixed generator. Pairwise samples are then created from correct/incorrect responses and used to train the model. This data generation approach cannot be adapted to create data for other tasks, like step-level evaluation, fundamentally limiting the task abilities of STE. A secondary concern is a lack of exposure to diverse response distributions; Evaluations in wang2024direct show that scaling training data with a simpler training recipe leads to better generalization across benchmarks compared to STE.
@@
Inclusion of direct judgment data. Past work (wang2024direct; cao2024compassjudger) has showed the importance of including direct judgment data samples to isolate judgment training signal. These are samples where the critique $c$ is omitted, and the input protocol $p$ is modified to prompt directly for a judgment. To precisely control the fraction of direct judgment data, we convert a fixed percent of $\mathcal{D}_{t}$ to direct judgment data by dropping generated critiques and modifying the input prompt accordingly. In App. D, we ablate the proportion of direct judgment data and show such data enables FARE to be prompted to exclude critiques for faster inference.
@@
Base models. We train two models starting from Qwen3-8B-Base (yang2025qwen3) and gpt-oss-20B (agarwal2025gpt), denoted FARE-8B and FARE-20B. We find Qwen3-8B (post-trained) to be over-trained, and therefore cold-start Qwen3-8B-Base from SFT data from Qwen2.5-32B-Instruct, which we denote Qwen3-8B-ColdStart. See § B.2 for additional details.
@@ 4.1 Core benchmarks
Setup. We evaluate along five diverse aspects: (ireasoning with JudgeBench (tan2024judgebench), ReasoningJudgeBench (RJB) (xu2025j4r), and PPE Correctness (frick2024evaluate), (iibias and robustness with RM-Bench (liu2024rm), (iiitool-use with When2Call (ross2025when2call), (ivstep-level error identification with ProcessBench (zheng2024processbench), and (vreference-based verification with VerifyBench (yan2025verifybench).
For RM-Bench and PPE Correctness pairwise benchmarks, we adopt default evaluation setups, running each benchmark once with a fixed random ordering of responses. For other pairwise benchmarks, we report consistent-accuracy (tan2024judgebench), where each test sample is run twice, swapping the order of response A and response B. If the evaluator selects a different response between runs (i.e., is positionally biased), then the sample is marked incorrect; if the evaluator is consistent, the judgment is graded against the ground-truth. For ProcessBench and VerifyBench, we report F1-score [^1] ProcessBench defines their reported F1-score differently from the traditional F1-score; See this link. and accuracy, respectively. We compare FARE against other finetuned generative and prompted evaluators. We report official numbers from past benchmarks, reporting sources in App. C. If necessary, we run each baseline using its own prompt template. For ProcessBench, we additionally compare against non-generative process reward models (PRMs).
@@
FARE also are extremely strong step-level evaluators, as shown by ProcessBench performance in Table 2. FARE-8B is the best small-sized generative critic model, outperforming the recently released StepWiser-7B (xiong2025stepwiser), a RL-trained specialized step-level evaluator, by 1.6 points. FARE-20B outperforms the specialized Qwen2.5-Math-72B-PRM by 6.1 points, even matching GPT-5. Most notably, FARE excel on the two most challenging splits, OlympiadBench and OmniMATH, with FARE-8B beating StepWiser-7B by 5.6 points and FARE-20B beating Qwen2.5-Math-72B-PRM by 7.3 points on average. Overall, FARE are not only capable outcome-level evaluators, but are able to find subtle mistakes that manifest at the step-level.
FARE are also capable verifiers, as shown in Table 3, with both models outperforming all reported baselines on VerifyBench-Hard (yan2025verifybench). In particular, FARE-20B excels beats the next best model, GPT-4o, by 12.5 absolute points. As we demonstrate in § 4.2, when used as verifiers during GRPO settings, FARE bring tangible benefits over typical string-matching verifiers.
Scaling inference-time compute. Recently, sampling parallel judgments and aggregating via majority vote, i.e., self-consistency (wang2022self), has been used to improve evaluator performance. In Table 4, we use self-consistency with 32 responses (SC@32) on PPE, comparing with J1 and DeepSeek-GRM (liu2025inference). DeepSeek-GRM also trains a MetaRM to perform judgment re-ranking at test-time. Across most splits, using SC@32 improves performance, with up to a 3.3 point improvement for FARE-20B on the MMLU-Pro split. Even without SC@32, FARE-8B beats DeepSeek-GRM-27B + MetaRM, with the gap widening with extra compute. Similarly, the gap between FARE-20B and J1-70B grows with extra compute. Interestingly, we see that across both our models, performance on MBPP+ slightly degrades, indicating that SC may not be the optimal way to use compute across all domains. Nonetheless, we observe gains in the aggregate.